OpenStudy (anonymous):

f(x)=2x^4x Logarithmic Derivatives

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

4 is the exponent? or 4x is?

OpenStudy (anonymous):

4x, 2^(4x)

jimthompson5910 (jim_thompson5910):

ok thx

OpenStudy (anonymous):

it says f(3)'

jimthompson5910 (jim_thompson5910):

y = 2x^(4x) ln(y) = ln( 2x^(4x) ) ln(y) = 4x*ln( 2x ) derive both sides with respect to x, then solve for y'

OpenStudy (anonymous):

1/y?

jimthompson5910 (jim_thompson5910):

the derivative of ln(y) with respect to x is ( y' )/y

OpenStudy (anonymous):

I see. is this for left hand?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

ok,

OpenStudy (anonymous):

what do i do for right hand?

jimthompson5910 (jim_thompson5910):

use the product rule

OpenStudy (anonymous):

4x(1/2x)+(in2x)(4)?

jimthompson5910 (jim_thompson5910):

ln is lower case L, then n not i then n

jimthompson5910 (jim_thompson5910):

oh wait, made a typo, hold on while I fix it

jimthompson5910 (jim_thompson5910):

y = 2x^(4x) ln(y) = ln( 2x^(4x) ) ln(y) = ln( 2) + ln( x^(4x) ) ln(y) = ln( x^(4x) ) + ln( 2) ln(y) = 4x*ln( x ) + ln( 2) ( y' )/y = 4*ln( x )+(4x/2x)*2 ( y' )/y = 4*ln( x )+4 y' = y ( 4*ln( x )+4 ) y' = 2x^(4x) * ( 4*ln( x )+4 ) I'll let you finish

OpenStudy (anonymous):

sure,

OpenStudy (anonymous):

mm product rule?

jimthompson5910 (jim_thompson5910):

so f ' (x) = 2x^(4x) * ( 4*ln( x )+4 ) and you plug in x = 3 to find f ' (3)

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

It says incorrect again

jimthompson5910 (jim_thompson5910):

what did you get

OpenStudy (anonymous):

i got 20777773655.1287321453297664

jimthompson5910 (jim_thompson5910):

way too big lol

OpenStudy (anonymous):

hahaha

jimthompson5910 (jim_thompson5910):

although, the answer is quite large...

jimthompson5910 (jim_thompson5910):

y = 2x^(4x) ln(y) = ln( 2x^(4x) ) ln(y) = ln( 2) + ln( x^(4x) ) ln(y) = ln( x^(4x) ) + ln( 2) ln(y) = 4x*ln( x ) + ln( 2) ( y' )/y = 4*ln( x )+(4x/2x)*2 ( y' )/y = 4*ln( x )+4 y' = y ( 4*ln( x )+4 ) y' = 2x^(4x) * ( 4*ln( x )+4 ) f ' (x) = 2x^(4x) * ( 4*ln( x )+4 ) f ' (3) = 2(3)^(4(3)) * ( 4*ln( 3 )+4 ) f ' (3) = 8,922,308.90641656

jimthompson5910 (jim_thompson5910):

of course, you'll leave off the commas they're just there for display purposes

OpenStudy (anonymous):

2.0777?

jimthompson5910 (jim_thompson5910):

how did you get that?

OpenStudy (anonymous):

by my cauclator,

OpenStudy (anonymous):

it says 2.077777366e10

jimthompson5910 (jim_thompson5910):

oh, that's close I guess, but what I got should be the correct answer

OpenStudy (anonymous):

what is ur answer?

jimthompson5910 (jim_thompson5910):

i posted it above

jimthompson5910 (jim_thompson5910):

see it?

OpenStudy (anonymous):

yes, omg, you are math talent

jimthompson5910 (jim_thompson5910):

no, that would be the calculator lol

jimthompson5910 (jim_thompson5910):

i wish i could calculate that in my head

OpenStudy (anonymous):

thank you, the way you write is clear to see, easy to learn

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

can i post more questions?

jimthompson5910 (jim_thompson5910):

sure